OpenStudy (anonymous):

Then graph the equation and use the graph to determine the cost of the car rental if the number of miles driven is 12.

OpenStudy (anonymous):

To graph, you need 2 points, find the cost when the mile driven is 0 and find the cost when the mile drive is 12.,

OpenStudy (anonymous):

Your cost equation is c=m/2+10 at m=0, c=0/2+10 = 10 at m=12, c=12/2 + 10=16

OpenStudy (anonymous):

Thank you oojih.com for your respose, now where do I place the points in the graph?

OpenStudy (anonymous):

you cost is a function of m; so like y is a function of x. Your c is a y-axis and m is an x-axis. Hope that helps

OpenStudy (anonymous):

Thank you again oojih.com your answer is very helpful.

OpenStudy (anonymous):

yw; let me know if it is still not clear

OpenStudy (anonymous):

will the points go on the $10 or the $16 y-axis and the x-axis point go on either the 12 or the 6?

OpenStudy (anonymous):

so, it is similar to plot y=x/2 + 10 2 points (0,10) and (12,16). Does this help?